OpenStudy (anonymous):

give the vertex and x-intercepts of f(x) -x squared +3x +2

OpenStudy (anonymous):

is this f(x) = x^2 +3x + 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so to find the x intercepts we have to factor the quadratic. so if we factor the quadratic we will get the following: (x+2)(x+1) therefore x = -1, -2. so they are your x intercepts. now to find the vertex the function will be at minimum. so we will have ... to find the min or max of a function you have to take the derivative. so derivative of this quad is 2x + 3 = 0 so 2x = -3 and x = -3/2 Hence the vertex of this is (-3/2, -2) i believe.

myininaya (myininaya):

or you can write the parabola in the form f(x)=a(x-h)^2+k

myininaya (myininaya):

vertex=(h,k)

myininaya (myininaya):

You need to know how to complete the square though

myininaya (myininaya):

so f(x)=x^2+3x+(3/2)^2+2+(3/2)^2=(x+3/2)^2+2+9/4=(x+3/2)^2+8/4+9/4=(x+3/2)^2+17/4 so the vertex is (-3/2,17/4)

myininaya (myininaya):

oops but i did something wrong

OpenStudy (anonymous):

I was about to say myin, but you are more intelligent than I. Fix it nao! :)

myininaya (myininaya):

f(x)=x^2+3x+(3/2)^2+2-(3/2)^2=(x+3/2)^2+2-9/4=(x+3/2)^2+8/4-9/4=(x+3/2)^2-1/4

myininaya (myininaya):

so the vertex is (-3/2,1/4) if I didn't screw up again

myininaya (myininaya):

darn it (-3/2,-1/4)

myininaya (myininaya):

and laxad found the x-intercepts above